# How the Schwarzschild–de Sitter horizons remain in thermal equilibrium at vastly different temperatures

### v1

@@ I Introduction
The metric is a solution of the vacuum Einstein equation with cosmological constant $\Lambda$ and can have two horizons, according to the values of $m$ and $\Lambda$. Here, we assume that the black hole radius is much smaller than the cosmological horizon size, i.e., $2m\ll H^{-1}=\sqrt{3/\Lambda}$. In this approximation, ${\cal T}_{\mathrm{H}}\approx{1}/(8\pi m)\gg{\cal T}_{\mathrm{GH}}\approx{H}/(2\pi)$. Then the temperatures of the two horizons are vastly different, with the black hole horizon being much hotter than the de Sitter one. It is often stated in the literature that, because of this difference, there cannot be thermal equilibrium in SdS space, except for the degenerate Nariai case [4] in which the two horizons coincide (which is not physically interesting); that multiple temperatures must be introduced; that the area law or other fundamental relations of black hole thermodynamics must be modified (e.g., [5, 6, 7, 8, 9, 10, 11, 12]). These conflicting proposals point to an incomplete understanding of multi-horizon spacetimes in general, and of SdS in particular. Here we propose an alternative view in which the two horizons remain in thermal equilibrium although their temperatures are different. Thermal equilibrium (defined as the time-independence of temperature and heat flux in the static frame) is achieved by letting the temperature vary with the radial position, interpolating between the Hawking and the Gibbons–Hawking temperatures (we provide explicitly the analytical temperature profile ${\cal T}(r)$ between the two horizons in the approximation of negligible back-reaction). Here by “temperature” we mean the temperature of the blackbody Hawking radiation between the two horizons, which will be treated as a radiation fluid at rest in the locally static geometry between these horizons. Since backreaction is neglected in the calculation of both the Hawking and the Gibbons–Hawking temperatures [1, 3], the test fluid approximation for the blackbody radiation is legitimate in the “usual” regime of moderate emission. The temperature profile is derived from a generalization of the Tolman–Ehrenfest criterion [13, 14, 15] for the thermal equilibrium of a test fluid at rest in a static spacetime. This generalized criterion applies in the presence of heat sources, which are absent in the “standard” situation conceived by Tolman.
@@ III Tolman–Ehrenfest criterion generalized to stationary heat conduction
The radial heat flux must be outgoing, which implies that $q_{0}>0$.
@@
${\cal T}(r)=\frac{1}{\sqrt{1-\frac{2m}{r}-\frac{\Lambda}{3}r^{2}}}\Bigg{[}{\cal T}_{0}+\frac{q_{0}}{2m{\cal K}}\Bigg{(}\ln x-\,\sum_{i}\,\frac{\left(\Lambda R_{i}^{2}-3\right)\ln\left(x-x_{i}\right)}{3\left(\Lambda R_{i}^{2}-1\right)}\Bigg{)}\Bigg{]}$ (3.18)
@@ IV Temperature regularization at the horizons
The divergence of the Tolman–Ehrenfest temperature at the black hole horizon and its relation with the Hawking temperature are well known in standard horizon thermodynamics [41, 42, 34]. Let us consider the Schwarzschild black hole: singularities are not new in thermodynamics and are related to the vacuum state of the quantum field used to compute the temperature (for example, the Boulware vacuum is singular on the black hole horizon). The Hawking temperature seen by a detector at constant radius $r$ is ${\cal T}=\kappa/\left(2\pi\right)$, where $\kappa$ is the surface gravity. To remain at fixed radius, this detector is accelerated with uniform acceleration $a$ and, in the inertial vacuum state, detects blackbody radiation at the Unruh temperature ${\cal T}_{\mathrm{U}}=a/\left(2\pi\right)$. In the Unruh vacuum the modes that, according to this observer, have propagated from the region near the black hole are populated (unlike the modes propagating from spatial infinity, which are not), at the temperature ${\cal T}=\kappa/\left(2\pi\sqrt{-k^{a}k_{a}}\right)$, where $k^{a}$ is the local timelike Killing vector. As the Schwarzschild horizon is approached, in the limit $r\to r_{\mathrm{S}}=2m$, the surface gravity is obtained as [2, 41]
@@
${\cal T}=\frac{a}{2\pi}\to\frac{\kappa}{2\pi\sqrt{-k^{c}k_{c}}}$ (4.2)
@@
${\cal T}\sqrt{-g_{00}}\simeq\frac{a}{2\pi}\simeq\frac{\kappa}{2\pi}$ (4.3)
@@
There is an even more important aspect regarding the apparent divergence of the temperature ${\cal T}$. There is increasing consensus that the Hawking radiation does not arise at the horizon but in a quantum atmosphere surrounding it, which is macroscopic [34, 35, 36, 37, 38, 39]. That is, this atmosphere is not confined to a Planck-sized near the horizon, but it extends to radii $r\simeq\frac{3\sqrt{3}}{4}\,r_{\mathrm{S}}\simeq 1.3r_{\mathrm{S}}$, as indicated by the fact that the Hawking flux indicates an emission area larger than the horizon area [34, 35, 36, 37, 38, 39]. Likewise, the wavelength of the typical Hawking quantum is not $r_{\mathrm{S}}$, but $\simeq 79r_{\mathrm{S}}$ [41, 34]. Presumably, similar considerations hold for the Gibbons–Hawking radiation associated with the de Sitter horizon.
@@ V Conclusions
Thermal equilibrium (or the lack thereof) between the two horizons of the SdS spacetime exhibits some similarities with heat conduction in a rod of finite length whose ends are kept at constant, different, temperatures. The temperature gradient maintains a steady heat flow and there is thermal equilibrium in the sense that both the heat flux density $\vec{q}$ and the linear temperature profile ${\cal T}(x)$ are time-independent. In the SdS case, it is the blackbody radiation between the two horizons (a perfect fluid) that is in thermal equilibrium with a radial temperature profile ${\cal T}(r)$ and the heat flux density (which is purely spatial in Eckart’s description. [^3] This feature, which implies superluminal heat flow and is clearly unrealistic, is corrected in a more refined causal thermodynamics formalisms [43, 44, 45, 46, 47, 48, 49], but the Eckart model is still the simplest and widely adopted model of dissipative fluid in relativity [50, 51].

### v2

@@ I Introduction
The metric is a solution of the vacuum Einstein equation with cosmological constant $\Lambda$ and can have two horizons, according to the values of $m$ and $\Lambda$. Here, we assume that the black hole radius is much smaller than the cosmological horizon size, i.e., $2m\ll H^{-1}=\sqrt{3/\Lambda}$. In this approximation, ${\cal T}_{\mathrm{H}}\approx{1}/(8\pi m)\gg{\cal T}_{\mathrm{GH}}\approx{H}/(2\pi)$. Then the temperatures of the two horizons are vastly different, with the black hole horizon being much hotter than the de Sitter one. It is often stated in the literature that, because of the temperature difference between the black hole and cosmological horizons, there cannot be thermal equilibrium in Schwarzschild-de Sitter (SdS) space—except for the degenerate Nariai case [4], where the two horizons coincide (a scenario which is not physically interesting). Consequently, multiple temperatures must be introduced, and some authors argue that the area law or other fundamental relations of black hole thermodynamics must be modified (e.g., [5, 6, 7, 8, 9, 10, 11, 12]). These conflicting proposals point to an incomplete understanding of multi-horizon spacetimes in general, and of SdS in particular. Here we propose an alternative view in which the two horizons remain in thermal equilibrium although their temperatures are different. Thermal equilibrium (defined as the time-independence of temperature and heat flux in the static frame) is achieved by letting the temperature vary with the radial position, interpolating between the Hawking and the Gibbons–Hawking temperatures (we provide explicitly the analytical temperature profile ${\cal T}(r)$ between the two horizons in the approximation of negligible back-reaction). Here by “temperature” we mean the temperature of the blackbody Hawking radiation between the two horizons, which will be treated as a radiation fluid at rest in the locally static geometry between these horizons. Since backreaction is neglected in the calculation of both the Hawking and the Gibbons–Hawking temperatures [1, 3], the test fluid approximation for the blackbody radiation is legitimate in the “usual” regime of moderate emission. The temperature profile is derived from a generalization of the Tolman–Ehrenfest criterion [13, 14, 15] for the thermal equilibrium of a test fluid at rest in a static spacetime. This generalized criterion applies in the presence of heat sources, which are absent in the “standard” situation conceived by Tolman.
@@ III Tolman–Ehrenfest criterion generalized to stationary heat conduction
$q^{r}$ diverges on the horizons due to the adopted coordinate system. However, the Tolman criterion is necessarily formulated in this coordinate system, which is adapted to the time symmetry. The radial heat flux must be outgoing, which implies that $q_{0}>0$.
@@
${\cal T}(r)=\frac{1}{\sqrt{1-\frac{2m}{r}-\frac{\Lambda}{3}r^{2}}}\Bigg{[}{\cal T}_{0}+\frac{q_{0}}{2m{\cal K}}\Bigg{(}\ln x-\,\sum_{i}\,\frac{\left(\Lambda R_{i}^{2}-3\right)\ln\left(x-x_{i}\right)}{3\left(\Lambda R_{i}^{2}-1\right)}\Bigg{)}\Bigg{]}\,,$ (3.18)
@@ IV Temperature regularization at the horizons
The divergence of the Tolman–Ehrenfest temperature at the black hole horizon and its relation with the Hawking temperature are well known in standard horizon thermodynamics [42, 43, 35]. Let us consider the Schwarzschild black hole: divergences are not new in thermodynamics and are related to the vacuum state of the quantum field used to compute the temperature (for example, the Boulware vacuum is singular on the black hole horizon). The Hawking temperature is ${\cal T}_{\rm H}=\kappa/\left(2\pi\right)$, where $\kappa$ is the surface gravity, [^3] The surface gravity $\kappa$ is defined globally [2] instead of locally or quasilocally. and it represents the temperature seen by a static distant observer ($r\to\infty$). To remain at fixed radius, a detector is accelerated with uniform acceleration $a$ and, in the inertial vacuum state, detects blackbody radiation at the Unruh temperature ${\cal T}_{\mathrm{U}}=a/\left(2\pi\right)$. In the Unruh vacuum the modes that, according to this observer, have propagated from the region near the black hole are populated (unlike the modes propagating from spatial infinity, which are not), at the temperature ${\cal T}=\kappa/\left(2\pi\sqrt{-k^{a}k_{a}}\right)$, where $k^{a}$ is the local timelike Killing vector. As the Schwarzschild horizon is approached, in the limit $r\to r_{\mathrm{S}}=2m$, the surface gravity is obtained as [2, 42]
@@
${\cal T}=\frac{a}{2\pi}\to\frac{\kappa}{2\pi\sqrt{-k^{c}k_{c}}}\,,$ (4.2)
@@
${\cal T}\sqrt{-g_{00}}\simeq\frac{a}{2\pi}{\sqrt{-g_{00}}}\simeq\frac{\kappa}{2\pi}\,$ (4.3)
@@
There is an even more important aspect regarding the apparent divergence of the temperature ${\cal T}$. There is increasing consensus that the Hawking radiation does not arise at the horizon but in a quantum atmosphere surrounding it, which is macroscopic [35, 36, 37, 38, 39, 40]. That is, this atmosphere is not confined to a Planck-sized near the horizon, but it extends to radii $r\simeq\frac{3\sqrt{3}}{4}\,r_{\mathrm{S}}\simeq 1.3r_{\mathrm{S}}$, as indicated by the fact that the Hawking flux indicates an emission area larger than the horizon area [34, 35, 36, 37, 38, 39, 40]. Likewise, the wavelength of the typical Hawking quantum is not $r_{\mathrm{S}}$, but $\simeq 79r_{\mathrm{S}}$ [42, 35]. Presumably, similar considerations hold for the Gibbons–Hawking radiation associated with the de Sitter horizon.
@@ V Conclusions
Thermal equilibrium (or the lack thereof) between the two horizons of the SdS spacetime exhibits some similarities with heat conduction in a rod of finite length whose ends are kept at constant, different, temperatures. The temperature gradient maintains a steady heat flow and there is thermal equilibrium in the sense that both the heat flux density $\vec{q}$ and the linear temperature profile ${\cal T}(x)$ are time-independent. In the SdS case, it is the blackbody radiation between the two horizons (a perfect fluid) that is in thermal equilibrium with a radial temperature profile ${\cal T}(r)$ and the heat flux density (which is purely spatial in Eckart’s description). This feature, which implies superluminal heat flow and is clearly unrealistic, is corrected in more refined causal thermodynamics formalisms [44, 45, 46, 47, 48, 49, 50], but the Eckart model is still the simplest and widely adopted model of dissipative fluid in relativity [51, 52].